OpenStudy (anonymous):

find the area of the region under the curve y=7-x^2 and above the x axis

OpenStudy (freckles):

first we need to find the intersections of y=7-x^2 and the x-axis (also called y=0) So first solve 7-x^2=0

OpenStudy (anonymous):

x= +- radical 7

OpenStudy (freckles):

 parabola looks something like this